OpenStudy (anonymous):

Solve the system by graphing. -5x-y=-11 2x-y=10

OpenStudy (anonymous):

@Algebraic! can you help me?

OpenStudy (anonymous):

Do you know how you can graph a straight line?

OpenStudy (anonymous):

Because this is the only thing that you have to do here. Graph 2 lines, find the intersection and abracadabra you are done.

OpenStudy (anonymous):

I can't get them into slope intercept form....

OpenStudy (anonymous):

-5x-y=-11 2x-y=10 You want them in y= somthing form For the first one if you add y to both sides and 11 you get -5x-y+y+11=-11+y+11 that is -5x+11=y or y=-5x+11

OpenStudy (anonymous):

Also for graphing you only need to find 2 points and connect them. -5x-y=-11 If x=0 -y=-11 so y=11 (0,11) is on the line If y=0 -5x=-11 so x=11/5 (11/5,0) is on the line too. Plot these two points and connect them so you have your line. Same with the other equation

OpenStudy (anonymous):

This was no help at all....

OpenStudy (anonymous):

I showed you 2 ways of doing this problem.

OpenStudy (anonymous):

I have to tell where the two lines intersect and they don't at all..

OpenStudy (anonymous):

Well they should, if they dont intersect than they are parallel, but these two are not.

OpenStudy (anonymous):

Where do they intersect? Because I did the problem three times and I can't see where they did..

OpenStudy (anonymous):

a, -5x-y=-11 b, 2x-y=10 I already expressed a, that was y=-5x+11 b, is y=2x-10 At x=0 a, takes y=11 b, takes y=-10 Now a, has a slope of -5 and b, has a slope of +2. So you can see that they have to meet at one point

OpenStudy (anonymous):

If you solve these equations you get the intersection. It will be at x=3, y=-4